OpenStudy (anonymous):

Solve the following system of equations. x + y = 3 2x + 2y = 6 (0, 3) (1.5, 1.5) no solution infinite number of solutions

OpenStudy (amoodarya):

they are the same if you divide second by 2 you will see

OpenStudy (amoodarya):

multiply 1st by -2 then add eq.

OpenStudy (smored):

infinite # of solutions

OpenStudy (anonymous):

In any system of equations, always use one equation to solve for one of the variables. I will use the first equation to solve for y in terms of x.\[x+y=3\]\[y=3-x\]Then substitute this new expression (3-x) for y in the other equation.\[2x+2(3-x)=6\]\[2x+6-2x=6\]\[2x-2x=6-6\]\[0=0\]When is 0 equal to 0? Always! So no matter what x and y are, they work. Therefore, there an infinite number of solutions.